ABSOLUTE INTENSITIES OF O_{3} LINES IN THE 9-11 μm REGION

\underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences Division, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; V. MALATHY DEVI, and D. CHRIS BENNER, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795.

We have completed our analysis of high-resolution (0.0027 cm^{−1}) absorption spectra of ozone^{a} to determine absolute intensities of 376 ^{16}O_{3} lines in the 9-11 μm region. The spectra were recorded at room temperature using the Fourier transform spectrometer at the McMath-Pierce facility of the National Solar Observatory at Kitt Peak. The ozone samples were contained in a glass cell having crossed IR-transmitting and UV-transmitting paths approximately 10 cm in each direction. The ozone partial pressures (0.3 to 0.5 Torr) in the cell were determined from measurements of the 254 nm UVabsorption, using the absorption cross-section of Mauersberger et al.^{b} as a reference standard. Only spectra for which the ozone partial pressure varied by < 1.0 % during the recording time were selected for analysis. A multispectrum nonlinear least-squares procedure^{c} was used to fit four spectra simultaneously to determine intensities for 366 lines in the P, Q, and R branches of the ν_{3} fundamental band and 10 lines in the R branch of the ν_{1} band. The absolute accuracy of these intensity values ranges from 2 % for the strongest, most well-determined lines to 4 or 5 % for the weakest lines measured. On average, our measured intensities are approximately 1 % larger than the values on the current HITRAN compilation.^{d} Our measurement set includes 44 ν_{3} lines in common with other recent experimental studies.^{e} ^{f} ^{g} Comparison of these various measurements shows excellent agreement for a few lines and adequate agreement (considering all possible sources of uncertainty and systematic errors) for the others.

aM. A. H. Smith, C. P. Rinsland, V. Malathy Devi, and D. Chris Benner, 52nd International Symposium on Molecular Spectroscopy, Paper RF03 (1997).

bK. Mauersberger, D. Hanson, J. Barnes, and J. Morton, J. Geophys. Res. \underline{92}, 8480-8482 (1987)

cD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith, and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

dL. S. Rothman et al., JQSRT \underline{48}, 469-507 (1992); L. S. Rothman et al., JQSRT, in press (1998).

eH. M. Pickett, D. B. Peterson, and J. S. Margolis, J. Geophys. Res. \underline{97}, 20,787-20,793 (1992).

fM. R. DeBacker, B. Parvitte, V. Zeninari, and D. Courtois, JQSRT \underline{54}, 1009-1018 (1995).

gM. R. DeBacker-Barilly and D. Courtois, Appl. Phys. B \underline{64}, 607-611 (1997).